Albuquerque, Inc., acquired 24,000 shares of Marmon Company several years ago for $820,000. At the acquisition date, Marmon reported a book value of $910,000, and Albuquerque assessed the fair value of the noncontrolling interest at $205,000. Any excess of acquisition-date fair value over book value was assigned to broadcast licenses with indefinite lives. Since the acquisition date and until this point, Marmon has issued no additional shares. No impairment has been recognized for the broadcast licenses. At the present time, Marmon reports $1,040,000 as total stockholders’ equity, which is broken down as follows:

Common stock ($12 par value) $ 360,000

Additional paid-in capital 450,000

Retained earnings 230,000

Total $ 1,040,000

View the following as independent situations: a. & b. Marmon sells 10,000 and 2,000 shares of previously unissued common stock to the public for $45 and $34 per share. Albuquerque purchased none of this stock. What journal entry should Albuquerque make to recognize the impact of this stock transaction? (If no entry is required for a transaction/event, select "No journal entry required" in the first account field. Do not round your intermediate calculations.)
Answer #1

Solution:

Acquisition interest at the beginning (24,000) / (360,000 / 12) = 80%

1.

Adjusted till date fair value ( 820,000 + 205,000) + (1,040,000 - 910,000) = 1,155,000

After the stock issue adjusted till date fair value (1,155,000) + (10,000 * 45) = 1,605,000

Current Interest after share issuance (24,000) / (30,000 + 10,000) = 60%

Investment equity balance before share issuance ( 820,000) + (1,040,000 - 910,000) * 80% = 924,000

Implied book value of investment after share issuance (1,605,000 * 60%) = 963,000

Increase in Additional paid-in Capital = 963,000 - 924,000 = 39,000

Journal entries:

Trans. Account Title and explanation Debit Credit 39,000 1 Invesetment in Marmon Additional paid in capital 39,000 (To record

Acquisition interest at the beginning (24,000) / (360,000 / 12) = 80%

2.

Adjusted till date fair value ( 820,000 + 205,000) + (1,040,000 - 910,000) = 1,155,000

After the stock issue adjusted till date fair value (1,155,000) + (2,000 * 34) = 1,223,000

Current Interest after share issuance (24,000) / (30,000 + 2,000) = 75%

Investment equity balance before share issuance ( 820,000) + (1,040,000 - 910,000) * 80% = 924,000

Implied book value of investment after share issuance (1,223,000 * 75%) = 917,250

Decrease in Additional paid-in Capital = 917,250 - 924,000 = -6,750

*** END ***

Do leave a thumbs up : )